# Predicting Drug-Gene Relations via Analogy Tasks with Word Embeddings

### v1

@@ 1.2.2 Pathway-wise setting by year
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}_{p}^{y}$ in Eq. (S7), we prepare the evaluation of the analogy tasks. For $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, using the projection operations in Eq. (10), we define $D_{p}^{y\mid I}:=\pi_{\mathcal{D}_{p}^{y}}(\mathcal{R}_{p}^{y\mid I})$ and $G_{p}^{y\mid I}:=\pi_{\mathcal{G}_{p}^{y}}(\mathcal{R}_{p}^{y\mid I})$. Also we define $D_{p}^{y}:=D_{p}^{y\mid(-\infty,\infty)}$ and $G_{p}^{y}:=G_{p}^{y\mid(-\infty,\infty)}$.
Similar to Eq. (11), for $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, we define $[d]_{p}^{y\mid I}\subset\mathcal{G}_{p}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D_{p}^{y\mid I}$, and $[g]_{p}^{y\mid I}\subset\mathcal{D}_{p}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G_{p}^{y\mid I}$. These are formally defined as follows:
@@
In particular, we define $[d]_{p}^{y}:=[d]_{p}^{y\mid(-\infty,\infty)}$ and $[g]_{p}^{y}:=[g]_{p}^{y\mid(-\infty,\infty)}$. Similar to settings P1, P2, Y1, and Y2, we perform the analogy tasks in the following four settings.
Setting P1Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes that are categorized in the same pathway $p$ as $d$ and whose relations appeared after year $y$ are considered correct. In other words, for a query drug $d\in D_{p}^{y\mid U_{y}}$, the set of answer genes is $[d]_{p}^{y\mid U_{y}}$. The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y\mid U_{y}}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}^{y\mid U_{y}}$, then the prediction is considered correct.
Setting P2Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct, regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not. In other words, for a query drug $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$, the set of answer genes is $[d]^{y\mid U_{y}}$. The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
@@
For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway as the drug $d$ are considered correct, regardless of whether their relations appeared up to or after year $y$. In other words, for a query drug $d\in D_{p}^{y}$, the set of answer genes is $[d]_{p}^{y}$. The search space is the set of all genes $\mathcal{G}^{y}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}^{y}$, then the prediction is considered correct.
@@
For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether they are categorized in the same pathway or not and whether their relations appeared up to or after year $y$. In other words, for a query drug $d\in\mathcal{D}_{p}^{y}\cap D^{y}$, the set of answer genes is $[d]^{y}$. The search space is the set of all genes $\mathcal{G}^{y}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct.
Note the following in settings P1Y1, P1Y2, P2Y1 and P2Y2:
  • In setting P1Y1, the search space is $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$. This ensures that predicted target genes, which are categorized in the same pathway $p$ as the drug $d$, have the relations that appeared only after year $y$.
  • In setting P2Y1, since $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}$ and $D_{p}^{y\mid U_{y}}=\pi_{\mathcal{D}_{p}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)=\pi_{\mathcal{D}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)\subset\pi_{\mathcal{D}^{y}}\left(\mathcal{R}^{y\mid U_{y}}\right)=D^{y\mid U_{y}}$, it follows that $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$. For $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$, there may be $g\in[d]^{y\mid U_{y}}\setminus[d]_{p}^{y\mid U_{y}}$. Therefore, in Eq. (S6), we consider $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ and $g\in[d]^{y\mid U_{y}}$, instead of $(d,g)\in\mathcal{R}_{p}^{y\mid U_{y}}$. The search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. This ensure that predicted target genes have the relations that appeared only after year $y$.
  • In setting P1Y2, for a query drug $d\in D_{p}^{y}$, to evaluate the ability to predict target genes categorized in the same pathway $p$ as the drug $d$, regardless of whether their relations appeared up to or after year $y$, we consider $(d,g)\in\mathcal{R}_{p}^{y}$ instead of $(d,g)\in\mathcal{R}_{p}^{y\mid U_{y}}$ in Eq. (20).
  • In setting P2Y2, since $D_{p}^{y}\subset\mathcal{D}_{p}^{y}$ and $D_{p}^{y}=\pi_{\mathcal{D}_{p}^{y}}\left(\mathcal{R}_{p}^{y}\right)=\pi_{\mathcal{D}^{y}}\left(\mathcal{R}_{p}^{y}\right)\subset\pi_{\mathcal{D}^{y}}\left(\mathcal{R}^{y}\right)=D^{y}$, it follows that $D_{p}^{y}\subset\mathcal{D}_{p}^{y}\cap D^{y}$. For $d\in\mathcal{D}_{p}^{y}\cap D^{y}$, there may be $g\in[d]^{y}\setminus[d]_{p}^{y}$. Therefore, in Eq. (S6), we consider $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ and $g\in[d]^{y}$, instead of $(d,g)\in\mathcal{R}_{p}^{y\mid U_{y}}$.
@@ G $d\in D$ $[d]$ $\mathcal{G}$
Y1 $d\in D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
Y2 $d\in D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
P1Y1 $d\in D_{p}^{y\mid U_{y}}$ $[d]_{p}^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$
P2Y1 $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
P1Y2 $d\in D_{p}^{y}$ $[d]_{p}^{y}$ $\mathcal{G}^{y}$
P2Y2 $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
@@
Table S2 shows the settings for predicting genes from drugs, as used in the main text. Similarly, Table S3 shows the settings for predicting drugs from genes. In Table S3, we adapted the notations used for predicting genes from drugs in Table S2 to those used for predicting drugs from genes, denoting them as ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$.
@@ 1.5.1 Conversion from MeSH ID to drug name
Embeddings Setting Method Centering Top1 Top10 MRR
Ours G Random 0.000 0.000 0.001
@@
Embed. Setting Method Top1 Top10 MRR
BCV ${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
@@
Ours ${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
@@
1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
$|\mathcal{D}^{y}|$ 2268 3275 5059 7125 9387 11748 14044 16827 19758 22717 25431
$|\mathcal{G}^{y}|$ 353 725 1782 3413 5976 10533 16949 24784 32264 39961 47509
$\left|\mathcal{R}^{y\mid L_{y}}\right|$ 12 37 117 256 499 859 1269 1637 1978 2354 2603
$\left|\mathcal{R}^{y\mid U_{y}}\right|$ 13 53 183 296 344 515 473 379 281 189 61
$\left|D^{y\mid U_{y}}\right|$ 12 48 154 272 288 369 325 256 189 131 48
$|D^{y}|$ 22 79 246 475 663 893 1047 1151 1269 1423 1495
$\left|G^{y\mid U_{y}}\right|$ 6 14 34 55 89 137 173 153 140 115 44
$|G^{y}|$ 8 20 46 74 122 215 325 398 477 540 570
$\mathrm{E}_{d\in D^{y\mid L_{y}}}\left\{\left|[d]^{y\mid L_{y}}\right|\right\}$ 1.091 1.028 1.104 1.133 1.188 1.372 1.500 1.613 1.665 1.722 1.760
$\mathrm{E}_{d\in D^{y\mid U_{y}}}\left\{\left|[d]^{y\mid U_{y}}\right|\right\}$ 1.083 1.104 1.188 1.088 1.194 1.396 1.455 1.480 1.487 1.443 1.271
$\mathrm{E}_{g\in G^{y\mid L_{y}}}\left\{\left|[g]^{y\mid L_{y}}\right|\right\}$ 3.000 2.467 4.179 5.020 5.940 5.335 4.789 4.547 4.425 4.467 4.591
$\mathrm{E}_{g\in G^{y\mid U_{y}}}\left\{\left|[g]^{y\mid U_{y}}\right|\right\}$ 2.167 3.786 5.382 5.382 3.865 3.759 2.734 2.477 2.007 1.643 1.386
Table S7: Statistics for settings Y1 and Y2.

### v2

@@ 1.2.4 Pathway-wise setting to predict unknown relations by year
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ in Eq. (S16), we prepare the evaluation of the analogy tasks. For $I\in\{L_{y},U_{y}\}$, using the projection operations in Eq. (10), we define $D_{p}^{y\mid I}:=\pi_{\mathcal{D}_{p}^{y}}(\mathcal{R}_{p}^{y\mid I})$ and $G_{p}^{y\mid I}:=\pi_{\mathcal{G}_{p}^{y}}(\mathcal{R}_{p}^{y\mid I})$.
Similar to Eq. (11), for $I\in\{L_{y},U_{y}\}$, we define $[d]_{p}^{y\mid I}\subset\mathcal{G}_{p}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D_{p}^{y\mid I}$, and $[g]_{p}^{y\mid I}\subset\mathcal{D}_{p}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G_{p}^{y\mid I}$. These are formally defined as follows:
@@
Similar to settings P1, P2, Y2, we perform the analogy tasks in the following four settings.
@@
For the target genes that have drug-gene relations with a drug $d$, only genes that are categorized in the same pathway $p$ as $d$ and whose relations appeared after year $y$ are considered correct. In other words, for a query drug $d\in D_{p}^{y\mid U_{y}}$, the set of answer genes is $[d]_{p}^{y\mid U_{y}}$. The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y\mid L_{y}},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y\mid U_{y}}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y\mid L_{y}},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}^{y\mid U_{y}}$, then the prediction is considered correct.
@@
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct, regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not. In other words, for a query drug $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$, the set of answer genes is $[d]^{y\mid U_{y}}$. The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y\mid L_{y}},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y\mid L_{y}},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
Note the following in settings P1Y2 and P2Y2:
  • In setting P1Y2, the search space is $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$. This ensures that predicted target genes, which are categorized in the same pathway $p$ as the drug $d$, have the relations that appeared only after year $y$.
  • In setting P2Y2, since $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}$ and $D_{p}^{y\mid U_{y}}=\pi_{\mathcal{D}_{p}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)=\pi_{\mathcal{D}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)\subset\pi_{\mathcal{D}^{y}}\left(\mathcal{R}^{y\mid U_{y}}\right)=D^{y\mid U_{y}}$, it follows that $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$. For $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$, there may be $g\in[d]^{y\mid U_{y}}\setminus[d]_{p}^{y\mid U_{y}}$. Therefore, in Eq. (S9), we consider $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ and $g\in[d]^{y\mid U_{y}}$, instead of $(d,g)\in\mathcal{R}_{p}^{y\mid U_{y}}$. The search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. This ensure that predicted target genes have the relations that appeared only after year $y$.
@@ G $d\in D$ $[d]$ $\mathcal{G}$
Y1 $d\in D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
Y2 $d\in D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
P1Y1 $d\in D_{p}^{y}$ $[d]_{p}^{y}$ $\mathcal{G}^{y}$
P2Y1 $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
P1Y2 $d\in D_{p}^{y\mid U_{y}}$ $[d]_{p}^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$
P2Y2 $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
@@
Model Version
GPT-3.5 gpt-3.5-turbo-0125
GPT-4 gpt-4-turbo-2024-04-09
GPT-4o gpt-4o-2024-05-13
Table S5: Version of each GPT model.
Role Prompt
system You are a biologist specializing in drug-target interactions. Your responses should be formatted as JSON, with data listed under the key ‘target_genes’.
user List the top $10$ potential target genes for the drug drug name.
Table S6: Prompt template used for the GPT series to predict the top 10 target genes for each query drug. Here, drug name is a specific name such as Bosutinib.
Table S2 shows the settings for predicting genes from drugs, as used in the main text. Similarly, Table S3 shows the settings for predicting drugs from genes. In Table S3, we adapted the notations used for predicting genes from drugs in Table S2 to those used for predicting drugs from genes, denoting them as ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$.
@@ 1.5.1 Conversion from MeSH ID to drug name
Model Setting Method Centering Top1 Top10 MRR
Our embeddings G Random 0.000 0.000 0.001
@@
Model Setting Method Top1 Top10 MRR
BioConceptVec ${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
@@
Our embeddings ${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
@@
Metric
Vocabulary Setting LLM Top1 Top10
BioConceptVec G GPT-3.5 0.611 0.785
GPT-4 0.665 0.795
GPT-4o 0.718 0.848
P1 GPT-3.5 0.581 0.780
GPT-4 0.633 0.792
GPT-4o 0.689 0.855
P2 GPT-3.5 0.627 0.803
GPT-4 0.685 0.813
GPT-4o 0.742 0.873
Our embeddings G GPT-3.5 0.637 0.804
GPT-4 0.693 0.818
GPT-4o 0.741 0.867
P1 GPT-3.5 0.604 0.796
GPT-4 0.654 0.808
GPT-4o 0.703 0.868
P2 GPT-3.5 0.653 0.821
GPT-4 0.710 0.832
GPT-4o 0.760 0.887
Table S9: Gene prediction performance for GPT models.
@@
1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
$|\mathcal{D}^{y}|$ 3275 5059 7125 9387 11748 14044 16827 19758 22717 25431 28284
$|\mathcal{G}^{y}|$ 725 1782 3413 5976 10533 16949 24784 32264 39961 47509 51057
$|\mathcal{R}^{y}|$ 128 429 807 1251 2337 3282 4054 4849 5551 5909 5968
$|D^{y}|$ 104 343 666 904 1253 1436 1566 1708 1875 1961 1980
$|G^{y}|$ 21 47 75 125 228 348 430 521 590 630 634
$\mathrm{E}_{d\in D^{y}}\{|[d]|\}$ 1.231 1.251 1.212 1.384 1.865 2.286 2.589 2.839 2.961 3.013 3.014
$\mathrm{E}_{g\in G^{y}}\{|[g]|\}$ 6.095 9.128 10.760 10.008 10.250 9.431 9.428 9.307 9.408 9.379 9.413
Table S10: Statistics for setting Y1.
Year
1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
$\left|\mathcal{R}^{y\mid L_{y}}\right|$ 37 117 256 499 859 1269 1637 1978 2354 2603
$\left|\mathcal{R}^{y\mid U_{y}}\right|$ 53 183 296 344 515 473 379 281 189 61
$\left|D^{y\mid U_{y}}\right|$ 48 154 272 288 369 325 256 189 131 48
$\left|G^{y\mid U_{y}}\right|$ 14 34 55 89 137 173 153 140 115 44
$\mathrm{E}_{d\in D^{y\mid L_{y}}}\left\{\left|[d]^{y\mid L_{y}}\right|\right\}$ 1.028 1.104 1.133 1.188 1.372 1.500 1.613 1.665 1.722 1.760
$\mathrm{E}_{d\in D^{y\mid U_{y}}}\left\{\left|[d]^{y\mid U_{y}}\right|\right\}$ 1.104 1.188 1.088 1.194 1.396 1.455 1.480 1.487 1.443 1.271
$\mathrm{E}_{g\in G^{y\mid L_{y}}}\left\{\left|[g]^{y\mid L_{y}}\right|\right\}$ 2.467 4.179 5.020 5.940 5.335 4.789 4.547 4.425 4.467 4.591
$\mathrm{E}_{g\in G^{y\mid U_{y}}}\left\{\left|[g]^{y\mid U_{y}}\right|\right\}$ 3.786 5.382 5.382 3.865 3.759 2.734 2.477 2.007 1.643 1.386
Table S11: Statistics for setting Y2.